Find the slope of the line that is (a) parallel and (b) perpendicular to the line through the pair of points.
(-7,-2) and (1,7)
The answer(s) has to be as an integer or simplified fraction

Answer by mananth(16946)

(-7,-2) and (1,7)
...
equation of a line passing through two points
y-y1= m(x-x1)
...
m= 7-(-2)/1-(-7)
m=9/8
...
y-(-2)=9/8(x-(-7))
y+2=9/8(x+7)
8(y+2)=9(x+7)
8y+16=9x+63
8y=9x+63-16
8y=9x+47
y = 9x/8 +(47/8)
slope of line parallel to this line = 9/8
slope of line perpendicular to this line = -8/9
...
m.ananth@hotmail.ca